# A Certified Unlearning Approach without Access to Source Data

### v2

@@ Appendix D Approximating Kullbeck-Leiber Distance
Also, by the Bretagnolle-Huber Inequality (Bretagnolle & Huber, 1979),
@@
Without direct access to exact samples from the target distribution, we approximate the input marginal KL divergence using energy-based modeling, as introduced in (Grathwohl et al., 2019). Energy-based models (EBMs) provide a flexible framework for modeling complex distributions by associating an energy score with each input, which corresponds to the unnormalized log-probability of the input under the target distribution. The normalized target distribution, denoted as $\rho(\boldsymbol{x})$, can be expressed as:
@@
By iteratively applying this update rule, the generated samples approximate the target distribution $\rho(\boldsymbol{x})$ without requiring explicit computation of the normalization constant $Z$. Given the collected samples from the distribution $\rho(\boldsymbol{x})$ along with the surrogate data samples, we then approximate the input marginal KL divergence using a Donsker-Varadhan variational representation (Donsker & Varadhan, 1983) of KL divergence.
@@ Appendix E Parameter Study and Implementation
We systematically evaluate our approach on both synthetic and real-world datasets to demonstrate its effectiveness in achieving certified unlearning. Unless stated otherwise, we use a linear training model with privacy parameters $\epsilon=5e^{3}$ and $\delta=1$, a forget ratio of $0.1$, and an $L_{2}$ regularization constant of $\lambda=0.01$. The loss function is assumed to be $\alpha$-strongly convex, $L$-Lipschitz, $\beta$-smooth, and $\gamma$-Hessian Lipschitz. Following prior works (Koh & Liang, 2017; Wu et al., 2023b, a; Zhang et al., 2024), we set $\alpha$, $L$, $\beta$, and $\gamma$ for each experimental setting as hyperparameters. We set $\alpha=1+\lambda$, $L=1$, $\beta=1$ and $\gamma=1$. Even if the added noise does not follows the exact theoretical requirements, it does not affect the theoretical soundness of the paper.
@@ Appendix F Additional Synthetic and Real-World Dataset Experiments
We conducted experiments on both synthetic and real datasets to justify the heuristic KL-divergence estimation (Section 4.2) and the corresponding empirical unlearning error $\hat{\Delta}$. In Figures 3 and 4, we plot both the “exact” and “approximated” results (our heuristic method) for the KL‐divergence and the respective noise $\sigma$. For the synthetic data experiments, the exact KL is computed using its closed-form for Gaussians. For real data experiments, since the exact KL divergence was not available, we estimated it using the Donsker-Varadhan bound as a reference, leveraging both exact and surrogate data samples. Figure 3, Figure 4, Tables 8 and 9 show our approximations closely match exact values.

### v3

@@ Appendix D Approximating Kullbeck-Leiber Distance
Also, by the Bretagnolle-Huber Inequality (bretagnolle1979),
@@
Without direct access to exact samples from the target distribution, we approximate the input marginal KL divergence using energy-based modeling, as introduced in (grathwohl_your_2019). Energy-based models (EBMs) provide a flexible framework for modeling complex distributions by associating an energy score with each input, which corresponds to the unnormalized log-probability of the input under the target distribution. The normalized target distribution, denoted as $\rho(\boldsymbol{x})$, can be expressed as:
@@
By iteratively applying this update rule, the generated samples approximate the target distribution $\rho(\boldsymbol{x})$ without requiring explicit computation of the normalization constant $Z$. Given the collected samples from the distribution $\rho(\boldsymbol{x})$ along with the surrogate data samples, we then approximate the input marginal KL divergence using a Donsker-Varadhan variational representation (donsker1983asymptotic) of KL divergence.
@@ Appendix E Parameter Study and Implementation
We systematically evaluate our approach on both synthetic and real-world datasets to demonstrate its effectiveness in achieving certified unlearning. Unless stated otherwise, we use a linear training model with privacy parameters $\epsilon=5e^{3}$ and $\delta=1$, a forget ratio of $0.1$, and an $L_{2}$ regularization constant of $\lambda=0.01$. The loss function is assumed to be $\alpha$-strongly convex, $L$-Lipschitz, $\beta$-smooth, and $\gamma$-Hessian Lipschitz. Following prior works (koh_understanding_2017; wue-edge-23; wu-graph-unlearning-23; zhang_towards_2024), we set $\alpha$, $L$, $\beta$, and $\gamma$ for each experimental setting as hyperparameters. We set $\alpha=1+\lambda$, $L=1$, $\beta=1$ and $\gamma=1$. Even if the added noise does not follows the exact theoretical requirements, it does not affect the theoretical soundness of the paper.
@@ Appendix F Additional Synthetic and Real-World Dataset Experiments
We conducted experiments on both synthetic and real datasets to justify the heuristic KL-divergence estimation (Section 4.2) and the corresponding empirical unlearning error $\hat{\Delta}$. In Figures 3 and 4, we plot both the “exact” and “approximated” results (our heuristic method) for the KL divergence and the respective noise $\sigma$. For the synthetic data experiments, the exact KL divergence is computed using its closed-form for Gaussians. For real data experiments, since the exact KL divergence was not available, we estimated it using the Donsker-Varadhan bound as a reference, leveraging both exact and surrogate data samples. Figure 3, Figure 4, Tables 8 and 9 show our approximations closely match exact values.